Injector Studies 

December 5th, 0800 - December 19th, 0000

 

Wednesday, 12/05/2012

Time
Description
Studiers
Comments

0745-0800

SR Fill  324 Singlets  Non Top-Up

OPS


0800-1600

LEUTL ACIS Valdiation


J. Forrestal et al

WR -31185
Requirements for the Validation:
1.  Both Synchrotron and LEUTL in ACIS Test Mode (Test Loads connected).
2.  Access to both Synchrotron and LEUTL tunnels.
3.  The Linac/PAR can stay in "Normal" Mode.

0800-0900

Linac Tunnel Access


OPS

John Oziemkowski  and Owen Obasohan (Tunnel Access Training)

0800-

Replace L1 klystron


A. Cours et al

WR - 31359

0900-1200

Linac Tunnel Access for PCGun water station re-piping.


R. Wright et al

WR -
31261


1200-1300

Restore all Linac / PAR for beam transport......

OPS


John Oziemkowski  and Owen Obasohan (Observe Tune-Up)


1300-1400

Test L1:RG1:QM4 power supply

T. Puttkammer


WR - 31305


1600-1900

Restore all injectors  for beam transport......

OPS




1900-2000


Injector Recovery / SR Fill


OPS


324 Singlets


2000-2400


L4/L5 RF Conditioning to 80MW


OPS


Record progress in operation on-line log

Please follow attached RF Conditioning.

Thursday, 12/06/2012

Time
Description
Studiers
Comment


0000 - 0700


L4/L5 RF Conditioning 80MW


OPS


Record progress in operation on-line log.


0700-0800


Injector Recovery/SR Fill


OPS 


324 Singlets


0800-0900

Linac Gallery Awareness with L3 modulator Start-up Demo.

S. Pasky

John Oziemkowski  and Owen Obasohan

0800-1400

Test booster current readback ADC card and interface and acquire waveforms. Also switch over to the new board and test the modification of operational software

CY Yao, Shifu Xu, Ju  Wang, H. Shang.


1400-1900

Restore all injectors for normal 375MeV beam transport......

OPS


John Oziemkowski  and Owen Obasohan (Observe Tune-Up)


1900-2000


Injector Recovery/SR Fill


OPS 


324 Singlets


2000-2400


Operator challenge


OPS


ACO
Discretion..........?
* Make a SCR save with your name in the file.
(Maintain RG2 current at 260ma)

Friday, 12/07/2012

Time
Description
Studiers
Comments

0000-0200

Restore all injectors for normal 375MeV beam transport......

OPS


Make and SCR save.


0200 - 0700


L4/L5 RF Conditioning to 90MW


OPS


Record progress in operation on-line log.
Please follow attached RF Conditioning


0700-0800


Injector Recovery/SR Fill


OPS 


324 Singlets
(Maintain RG2 current at 260ma) 


0800-1400

Test booster current readback ADC card and interface and acquire waveforms including optimizing bcontrol, autoRampCorrection and injection controllaw under the new board.

CY Yao, Shifu Xu, Ju  Wang


1400-1900

Restore all injectors for normal 375MeV beam transport......

OPS


John Oziemkowski  and Owen Obasohan (Observe Tune-Up)


1900-2000


Injector Recovery/SR Fill


OPS 


324 Singlets
(Maintain RG2 current at 260ma) 


2000 - 2400


L4/L5 RF Conditioning to 100MW


OPS


Record progress in operation on-line log.

Saturday, 12/08/2012

Time
Description
Studiers
Comments


0000 - 0700


L4/5 RF Conditioning to 100MW.


OPS


Record progress in operation on-line log.


0700-0800


Injector Recovery/SR Fill


OPS 


324 Singlets
(Maintain RG2 current at 260ma)


0800-1900


L4/5 RF Conditioning to 100MW.


OPS




1900-2000


Injector Recovery/SR Fill 


OPS


324 Singlets
(Maintain RG2 current at 260ma)


2000-2400


Slowly elevate and maintain 110 MW SLED power on both L4 and L5.


OPS


Sunday, 12/09/2012

Time
Description
Studiers
Comments


0000-0700


For conditioning purposes slowly elevate and maintain 110 MW SLED power on both L4 and L5.


OPS



Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets
(Maintain RG2 current at 260ma)


0800-1900


Slowly elevate and maintain 110 MW SLED power on both L4 and L5.


OPS



Adjust as needed for SR fills.


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets
(Maintain RG2 current at 260ma)


2000-2400


Slowly elevate and maintain 120 MW SLED power on both L4 and L5.


OPS


Monday, 12/10/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1400

Measure orbit response and emittance of the booster.

CY Yao


1400-1900

Restore all injectors for normal 375MeV beam transport......

OPS




1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Slowly elevate and maintain 90 MW SLED power on both L4 and L5.


OPS


Tuesday, 12/11/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0800 - 2400

Machine Studies Period

Wednesday, 12/12/2012

Time
Description
Studiers
Comments

0000- 0800

Machine Studies Period


0800-1400

Validate RG1 kicker


S. Pasky




0800-1400

BIK to look for the zero read back problem


L. Donley


WR- 31266


0800-1400

BIS mod to prevent charging supply from over charging


A. Hillman


WR- 31267


1400-1900

SetUp a 75Mev linac file for booster bypass pitching magnets radiation measurements.

N. Sereno et al

The LEUTL Tunnel and Booster must be in Beam Permit Mode followed by the Booster Bypass pitching magnets unlocked.

Administrative locks are #25 and #35 for BB-BM1 and BB-BM2 respectively.

Panel in building 412.
BB-BM1, TRP-J12 ckt 11&13 and BB-BM2 TRP-J12 ckt 15&17.

"The Administrative locks #25 and #35 for BB-BM1 and BB-BM2 must be restored prior to taking the LEUTL tunnel out of beam permit"!!!!


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Slowly elevate and maintain 90MW SLED power on both L4 and L5.


OPS



Thursday, 12/13/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1400

Validate RG2 kicker


S. Pasky




0800-1400

Investigate BIS gain values in hardware and software to match setpoint to actual output.


A. Hillman


WR- 31268


1400-1700

 Check LTP/PB/PTB and BB optics.

N.Sereno




1700-1900

SetUp linac 75MeV file for booster bypass pitching magnets radiation measurements.

N. Sereno, M. Borland, J. Vacca

The LEUTL Tunnel and Booster must be in Beam Permit Mode followed by the Booster Bypass pitching magnets unlocked.

Administrative locks are #25 and #35 for BB-BM1 and BB-BM2 respectively.

Panel in building 412.
BB-BM1, TRP-J12 ckt 11&13 and BB-BM2 TRP-J12 ckt 15&17.

"The Administrative locks #25 and #35 for BB-BM1 and BB-BM2 must be restored prior to taking the LEUTL tunnel out of beam permit"!!!!


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


For conditioning purposes elevate and maintain 110 MW SLED power on both L4 and L5.


OPS


Friday, 12/14/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1600


Test Interleaving ........


N. Sereno et al ..


Test interleaving using RG2 and LEUTL optics in place.
Goal would be to see if we can inject into the booster this way.


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Slowly elevate and maintain 90MW SLED power on both L4 and L5.


OPS


Saturday, 12/15/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1900


Slowly elevate and maintain 90MW SLED power on both L4 and L5.


OPS


Adjust as needed for SR fills.


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Slowly elevate and maintain 90MW SLED power on both L4 and L5.


OPS


Adjust as needed for SR fills.

Sunday, 12/16/2012

Time
Description
Studiers
Comments

0000-0700


Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1900


Slowly elevate and maintain 90MW SLED power on both L4 and L5.


OPS



Adjust as needed for SR fills.


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Maintain 90MW SLED power on both L4 and L5.


OPS


Adjust as needed for SR fills.

Monday, 12/17/2012

Time
Description
Studiers
Comments

0000-0700


Maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1400

Test booster current readback ADC card and interface and acquire waveforms

CY Yao, Shifu Xu, Ju  Wang


1400-1900


Measure booster emittance versus couplign with BTS:FS3 of low emittance lattice


CY Yao




1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Maintain 90MW SLED power on both L4 and L5.


OPS


Adjust as needed for SR fills.

Tuesday, 12/18/2012

Time
Description
Studiers
Comments

0000-0700

Slowly elevate and maintain 90MW SLED power on both L4 and L5. 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1200

Measure orbit response matrix data


CY Yao



1200-1900


COTR experiment


Sereno, Lumpkin, Dooling


RG2 500ma


1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets

2000-2400


Continued COTR experiment


Sereno, Lumpkin, Dooling


RG2 500ma

Wednesday, 12/19/2012

Time
Description
Studiers
Comments

0000-0700

Continued COTR experiment  followed by
90MW SLED power on both L4 and L5.
 


OPS


Adjust as needed for SR fills.


0700-0800


Injector Recovery/SR Fill. 


OPS


324 Singlets


0800-1300

Vertical corrector ramp tuning of 132-nm lattice


CY Yao




1300-1600

Booster emittance  measurement.


CY Yao




1900-2000


Injector Recovery/SR Fill.


OPS


324 Singlets


2000-2400


Prepare for Shutdown


OPS